OpenStudy (anonymous):

what is the quadratic formula how do you use it to solve x2 + 6x = –3.

OpenStudy (roadjester):

is that x^2+6x=-3?

OpenStudy (roadjester):

\[ax^2+bx+c=0\] is the general form you need in order to use the quadratic formula

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

How do you use the quadratic formula

OpenStudy (roadjester):

once you have it in that form, you use:\[x={-b \pm \sqrt{b^2-4ac} \over {2a}}\]

OpenStudy (roadjester):

That is the quadratic formula. Just substitute the values for a, b, and c

OpenStudy (roadjester):

Just remember that the function MUST equal 0

OpenStudy (anonymous):

Is it -6 -6 over /2

OpenStudy (roadjester):

so you have to move the -3 to the other side so that you have x^2 +6x+3=0

OpenStudy (roadjester):

You have two answers, the \[\pm\] so -6\[\pm\] sqrt(36-4(1)(3)) all divided by 2

OpenStudy (anonymous):

-3+- 6

OpenStudy (roadjester):

\[x={{-6\pm \sqrt{36-4(1)(3)}}\over {2(1)}}\]

OpenStudy (roadjester):

Do you know how to simplify that?